# On the volume of K-semistable Fano manifolds

### v1

@@ 4. Volume estimates and minimal rational curves
We use the notation $l_{X}$ from (6)-(7). If $l_{X}=n+1$, then as mentioned in Example 2.9 $X$ must be ${\mathbb{P}}^{n}$, and if $l_{X}=n$ then $X$ is either $Q^{n}$ or the blowup of ${\mathbb{P}}^{n}$ along a smooth subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ in a hyperplane. When $d_{Y}\geq 2$, by Lemma 4.2 below we know $(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}<2n^{n}$. When $d_{Y}=1$, by Lemma 4.3 below we know that ${\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is not a K-semistable Fano manifold, so it is excluded. Thus we verified the Theorem in the case $l_{X}\geq n$. When $l_{X}<n$, there are minimal rational curves of degree $d\in\{2,\dots,n-1\}$. For $12<d<n-$ 1, by Theorem 1.4 and Lemma 3.5, ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})<2n^{n}$. When $d=2$, by Theorem 3.1 we get ${\rm vol}(X)\leq 2n^{n}$ with equality if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
@@
Suppose $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\cdots,n\}$ that is contained in a hyperplane (See Example 2.9). Then
@@
where $s_{k-2}(N_{Y/{\mathbb{P}}^{n}})$ denote the $(k-2)$-th Segre class of normal bundle $N_{Y/{\mathbb{P}}^{n}}$. If $d=1$, then $\sum_{i=0}^{k-2}d^{i}=k-1$. By the combinatorial identity $\binom{n}{k}(k-1)=n\binom{n-1}{k-1}-\binom{n}{k}$,
@@
Suppose $\pi\colon X\rightarrow{\mathbb{P}}^{n}$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y\cong{\mathbb{P}}^{n-2}$ of degree $d_{Y}=1$ that is contained in a hyperplane, then $X$ is K-unstable Fano manifold.
@@
Let $E=\pi^{-1}({\mathbb{P}}^{n-2})$ be the exceptional divisor. And let $H=\pi^{*}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ be the pullback of a hyperplane section of ${\mathbb{P}}^{n}$. Note that the nef cone is $\text{Nef}(X)={\mathbb{R}}_{\geqslant 0}[H]+{\mathbb{R}}_{\geqslant 0}[H-E]$. Then the divisor $-K_{X}-tE=(n+1)H-(1+t)E$ is nef if and only if $0\leqslant t\leqslant n$. When $0\leqslant t\leqslant n$, by the similar computation as in Lemma 4.2,
@@
Note that $X={\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is a toric Fano variety, so one can also prove lemma 4.3 via toric geometry. The minimal generators of the fan $\Sigma\subseteq{\mathbb{R}}^{n}$ is given by $e_{1}=(1,0\cdots,0),\cdots,e_{n}=(0,\cdots,0,1),e_{0}=(-1,-1,\cdots,-1),f=(1,1,0,\cdots,0)$. Here the ray generated by $f$ corresponds to the exceptional divisor $E=\pi^{-1}({\mathbb{P}}^{n-2})$ by the cone-orbit correspondence. Then the moment polytope with respect to $-K_{X}$ is given by
@@
Let $i(X)\coloneqq{\rm max}\{r\in{\mathbb{N}}^{*}\mid-K_{X}\sim_{{\mathbb{Q}}}rL\,\text{for some ample line bundle }\,L\}$ be the Fano index of $X$. Then the above argument shows that ${\widehat{\rm vol}}({\mathcal{C}},o)\leqslant i(X)\cdot(-K_{X})^{n}$. We know that for smooth Fano manifold, the Fano index always satisfies $i(X)\leqslant n+1$ and $i(X)=n+1$ if and only if $X\cong{\mathbb{P}}^{n}$ and $i(X)=n$ if and only if $X\cong Q\subset{\mathbb{P}}^{n+1}$ is a smooth quadric hypersurface ([KO73]). If $X\cong{\mathbb{P}}^{n}$ and $r=i({\mathbb{P}}^{n})=n+1$, then $({\mathcal{C}}({\mathbb{P}}^{n},{\mathcal{O}}(1)),o)\cong({\mathbb{C}}^{n},0)$ is a smooth point, which is excluded in the ODP conjecture. If $r<i({\mathbb{P}}^{n})=n+1$, then there exists some $p\in{\mathbb{Z}}_{>1}$ such that $rp=i({\mathbb{P}}^{n})=n+1$. Then ${\widehat{\rm vol}}({\mathcal{C}}({\mathbb{P}}^{n},L),o)=r(-K_{{\mathbb{P}}^{n}})^{n}=(n+1)^{n+1}/p<2n^{n+1}$. So we can assume $X\ncong{\mathbb{P}}^{n}$. Then by Theorem 1.1, we have $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q$. Combined with Kobayashi-Ochiai’s result $i(X)\leqslant n$, we get
@@ 4.2. Examples and questions
Consider the ${\mathbb{C}}^{*}$ action on ${\mathbb{P}}^{n+1}$ with weights $(0,2,0,2,\underbrace{1,\dots,1}_{n-3})$ on the homogeneous coordinates. The ${\mathbb{C}}^{*}$ action preserves the quadric hypersurface and gives rise to a divisorial valuation whose center on $X$ is the line $C$. This divisorial valuation can also be obtained via a weighted blowup along the line $C$. Precisely, in the $\mathbf{u}$-coordinates, we perform the weighted blow along $C\cap U_{0}=\{u_{3}=u_{4}=\cdots=u_{n+1}=0\}$ with weights $(0,2,1,\dots,1)$. In the $\mathbf{v}$ coordinates, we also have the weights $(0,2,1,\dots,1)$. The weighted blowup $\mu:\hat{X}\rightarrow X$ in this example is globally defined and the exceptional divisor $E$ is isomorphis to a weighted projective ${\mathbb{P}}(2,\underbrace{1,\dots,1}_{n-2})$-bundle over ${\mathbb{P}}^{1}$. We can calculate the volume function ${\rm vol}(-K_{X}-xE)=n^{n}{\rm vol}(H-\xi E)$ where $\xi=\frac{x}{n}$. Interestingly the situation turns out to be different for $n=3$ and $n\geq 4$. When $n=3$, the nef threshold (Seshadri constant) of $H$ with respect to $E$ coincides with the pseudo-effective threshold and is equal to 2. We have the expression:
@@
where we used a weighted analogue of the usual intersection formula via Segre classes for projective bundles. It is not surprising that the expression of this piece coincides with the piece in the expression of (20) when $x=n\xi\leq n$. The second piece is more difficult to get. One way to get it is to use the fact that the measure $-\frac{1}{n!}d{\rm vol}(H-tE)$ is the Duistermaat-Heckman measure of the previously mentioned Hamiltonian $S^{1}$ action (see [BHJ17]) and then use the symmetry of its density function around $\xi=1$ in our case. We can also directly calculate the Duistermaat-Heckman measure of $S^{1}$-action by using a well-established localization/Fourier formula ([BGV92, 7.4]) and for the interest of the reader we will provide such a derivation in appendix.
@@
Let $X\subset{\mathbb{P}}^{n+1}$ be a Fano hypersurface of degree $1\leq b\leq n$. Then the minimal rational curve is of splitting type ${\mathcal{O}}(1)^{n-b}\oplus{\mathcal{O}}^{b-1}$ so that $d=n-b+2$ (see [Kol96, Exercise V.4.4]). The volume of $X$ is given by $V_{d}=d^{n}(n-d+2)=(n+2-b)^{n}\cdot b$. One can easily verify that $V_{d}$ is strictly smaller than ${\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ except when $b=1$ or $b=2$.
Question 4.8.
Is there a K-semistable Fano manifold $X$ with ${\rm vol}(X)={\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ that contains a minimal rational curve of degree $13\leq d\leq n-$ 1 but is not isomorphic to ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$?
Our results imply that when $d=n$, there is just one such $X$, i.e. the quadric hypersurface $Q^{n}$, while when $d=2$, there is no such $X$. Moreover, such an $X$ must satisfy $l_{X}=d$ (see (7)).
@@
If $a_{0}=\cdots=a_{n-1}=1,\,a_{n}=2,\,a_{n+1}=n+1$ and $d=2n+2$. Namely, $X_{d}$ is a general degree $22n+$ 2 hypersurface in ${\mathbb{P}}(1^{n},2,n+1)$. Then $(-K_{X_{2n+2}})^{n}=1$. By [IF00, Thm 8.1], we know that $X_{2n+2}$ is smooth when $n$ is even. By [ST24, Thm 1.3], we know $\delta(X_{2n+2})\geqslant(n+1)/2>1$, so $X_{2n+2}$ is a smooth K-stable Fano manifold with minimal anti-canonical volume $(-K_{X_{2n+2}})^{n}=1$ for all even $n$. When $n$ is odd, the general hypersurface $X_{2n+2}\subset{\mathbb{P}}(1^{n},2,n+1)$ is singular. (After communicating with K. Fujita, we find that this type of example already appears in [KSC04, Theorem 5.22].)
@@
Let $Y$ be the double cover of ${\mathbb{P}}^{n}$ ramified along a degree $2n$ smooth hypersurface $D$. Then by Hurwitz’s formula, $K_{Y}=\phi^{*}(K_{{\mathbb{P}}^{n}}+\frac{1}{2}D)$. So $(-K_{Y})^{n}=\deg(\phi)\cdot(-K_{{\mathbb{P}}^{n}}-\frac{1}{2}D)^{n}=2$. By the result of [LZ22, Zhu21], we know K-stability of $Y$ is equivalent to the K-stability of the log Fano pair $({\mathbb{P}}^{n},\frac{1}{2}D)$. Since we assume $D$ is a smooth hypersurface, then the log Fano pair $({\mathbb{P}}^{n},(\frac{n+1}{2n}-\varepsilon)D)$ is klt, so it is uniformly K-stable by [ADL23, Theorem 2.10] for $0<\varepsilon\ll 1$. Since $D\sim_{{\mathbb{Q}}}-K_{{\mathbb{P}}^{n}}$, by the interpolation of K-stability ([ADL24, Prop 2.13]), we conclude that $({\mathbb{P}}^{n},\frac{1}{2}D)$ is K-stable. Thus, $Y$ is a smooth K-stable Fano manifold with volume $(-K_{Y})^{n}=2$ for all dimensions $n\geqslant 2$. Note that $Y$ contains a degree 2 free-immersed rational curve with $n-1$ nodes (see [Kol96, IV.2.12.3]).

### v2

@@ 4.3. Minimal rational curves with degree $d\geq 3$
We use the notation $l_{X}$ from (6)-(7). If $l_{X}=n+1$, then by Theorem 2.9, $X$ must be ${\mathbb{P}}^{n}$, and if $l_{X}=n$ then $X$ is either $Q^{n}$ or the blowup of ${\mathbb{P}}^{n}$ along a smooth codimension two subvariety $Y$ of degree $d_{Y}\in\{1,\dots,n\}$ in a hyperplane. When $d_{Y}\geq 2$, by Lemma 4.12 below we know $(-K_{{\rm Bl}_{Y}{\mathbb{P}}^{n}})^{n}<2n^{n}$. When $d_{Y}=1$, by Lemma 4.13 below we know that ${\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is not a K-semistable Fano manifold, so it is excluded. Thus, we verified the Theorem 1.1 in the case $l_{X}\geq n$. When $l_{X}<n$, there are minimal rational curves of degree $d\in\{2,\dots,n-1\}$. For $3\leq d\leq n-1$, by Proposition 4.8, ${\rm vol}(X)<2n^{n}$. When $d=2$, by Theorem 1.4 and Proposition 4.2 we get ${\rm vol}(X)\leq 2n^{n}$ with equality if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$. ∎
@@
Suppose $X$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y$ of degree $d_{Y}\in\{1,\cdots,n\}$ that is contained in a hyperplane (See Theorem 2.9). Then
@@
where $s_{k-2}(N_{Y/{\mathbb{P}}^{n}})$ denote the (k2)(k-2)-th Segre class of normal bundle $N_{Y/{\mathbb{P}}^{n}}$. If $d=1$, then $\sum_{i=0}^{k-2}d^{i}=k-1$. By the combinatorial identity $\binom{n}{k}(k-1)=n\binom{n-1}{k-1}-\binom{n}{k}$,
@@
Suppose $\pi\colon X\rightarrow{\mathbb{P}}^{n}$ is the blowup of ${\mathbb{P}}^{n}$ along a subvariety $Y\cong{\mathbb{P}}^{n-2}$ of degree $d_{Y}=1$ contained in a hyperplane of ${\mathbb{P}}^{n}$, then $X$ is K-unstable.
@@
Let $E=\pi^{-1}({\mathbb{P}}^{n-2})$ be the exceptional divisor. And let $H=\pi^{*}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ be the pullback of a hyperplane section of ${\mathbb{P}}^{n}$. Note that the nef cone is $\text{Nef}(X)={\mathbb{R}}_{\geqslant 0}[H]+{\mathbb{R}}_{\geqslant 0}[H-E]$. Then the divisor $K_{X}-tE=(n+1)H-(1+t)E$ is nef if and only if $0\leqslant t\leqslant n$. When $0\leqslant t\leqslant n$, by the similar computation as in Lemma 4.12,
@@
Note that $X={\rm Bl}_{{\mathbb{P}}^{n-2}}{\mathbb{P}}^{n}$ is a toric Fano variety, so one can also prove lemma 4.13 via toric geometry. The minimal generators of the fan $\Sigma\subseteq{\mathbb{R}}^{n}$ is given by $e_{1}=(1,0\cdots,0),\cdots,e_{n}=(0,\cdots,0,1),e_{0}=(-1,-1,\cdots,-1),f=(1,1,0,\cdots,0)$. Here the ray generated by $f$ corresponds to the exceptional divisor $E=\pi^{-1}({\mathbb{P}}^{n-2})$ by the cone-orbit correspondence. Then the moment polytope with respect to $K_{X}$ is given by
@@
Let $i(X)\coloneqq{\rm max}\{r\in{\mathbb{N}}^{*}\mid-K_{X}\sim_{{\mathbb{Q}}}rL\,\text{for some ample line bundle }\,L\}$ be the Fano index of $X$. Then the above argument shows that ${\widehat{\rm vol}}({\mathcal{C}},o)\leqslant i(X)\cdot(-K_{X})^{n}$. We know that for smooth Fano manifold, the Fano index always satisfies $i(X)\leqslant n+1$ and $i(X)=n+1$ if and only if $X\cong{\mathbb{P}}^{n}$ and $i(X)=n$ if and only if $X\cong Q\subset{\mathbb{P}}^{n+1}$ is a smooth quadric hypersurface ([KO73]).
If $X\cong{\mathbb{P}}^{n}$ and $r=i({\mathbb{P}}^{n})=n+1$, then $({\mathcal{C}}({\mathbb{P}}^{n},{\mathcal{O}}(1)),o)\cong({\mathbb{C}}^{n},0)$ is a smooth point, which is excluded in the ODP conjecture. If $r<i({\mathbb{P}}^{n})=n+1$, then there exists some $p\in{\mathbb{Z}}_{>1}$ such that $rp=i({\mathbb{P}}^{n})=n+1$. Then ${\widehat{\rm vol}}({\mathcal{C}}({\mathbb{P}}^{n},L),o)=r(-K_{{\mathbb{P}}^{n}})^{n}=(n+1)^{n+1}/p<2n^{n+1}$. So we can assume $X\ncong{\mathbb{P}}^{n}$. Then by Theorem 1.1, we have $(-K_{X})^{n}\leqslant 2n^{n}$ and the equality holds if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q$. Combined with Kobayashi-Ochiai’s result $i(X)\leqslant n$, we get
@@ 4.5. Examples and questions
$v_{0}=\frac{Z_{2}}{Z_{2}}=u_{2}^{-1},v_{1}=\frac{u_{1}}{u_{2}}=-u_{3}-\sum_{k=4}^{n+1}\frac{u_{k}^{2}}{u_{2}},v_{k}=\frac{Z_{k}}{Z_{2}}=\frac{Z_{k}}{Z_{0}}\frac{Z_{0}}{Z_{2}}=u_{k}\cdot u_{2}^{-1}$
@@
Consider the $\mathbb{C}$$^{*}$ action on ${\mathbb{P}}^{n+1}$ with weights $(0,2,0,2,\underbrace{1,\dots,1}_{n-3})$ on the homogeneous coordinates. The $\mathbb{C}$$^{*}$ action preserves the quadric hypersurface and gives rise to a divisorial valuation whose center on $X$ is the line $C$. This divisorial valuation can also be obtained via a weighted blowup along the line $C$. Precisely, in the $\mathbf{u}$-coordinates, we perform the weighted blow along $C\cap U_{0}=\{u_{3}=u_{4}=\cdots=u_{n+1}=0\}$ with weights $(0,2,1,\dots,1)$. In the $\mathbf{v}$ coordinates, we also have the weights $(0,2,1,\dots,1)$. We see that the weighted blowup $\mu:\hat{X}\rightarrow X$ is indeed globally defined and the exceptional divisor $E$ is isomorphis to a weighted projective ${\mathbb{P}}(2,\underbrace{1,\dots,1}_{n-2})$-bundle over ${\mathbb{P}}^{1}$.
We can calculate the volume function ${\rm vol}(-K_{X}-xE)=n^{n}{\rm vol}(H-\xi E)$ where $\xi=\frac{x}{n}$. Interestingly the situation turns out to be different for $n=3$ and $n\geq 4$. When $n=3$, the nef threshold (Seshadri constant) of $H$ with respect to $E$ coincides with the pseudo-effective threshold and is equal to 2. We have the expression:
@@
where we used a weighted analogue of the usual intersection formula via Segre classes for projective bundles. It is not surprising that the expression of this piece coincides with the piece in the expression of (23) when $x=n\xi\leq n$.
The second piece is more difficult to get. One way to get it is to use the fact that the measure $\frac{1}{n!}d{\rm vol}(H-tE)$ is the Duistermaat-Heckman measure of the previously mentioned Hamiltonian $S^{1}$ action (see [BHJ17]) and then use the symmetry of its density function around $\xi=1$ in our case. We can also directly calculate the Duistermaat-Heckman measure of $S^{1}$-action by using a well-established localization/Fourier formula ([BGV92, 7.4]) and for the interest of the reader we will provide such a derivation in Appendix B. It is straightforward to verify the estimate ${\rm vol}(-K_{X}-xE)\geq V-\phi(x)$ (see (30)) for this example which is reduced to the inequality: for $n\leq x\leq 2n$,
$F(x):=\frac{n^{2}}{2}(2n-x)^{n-1}-\frac{n-2}{2}(2n-x)^{n}+\frac{1}{2}n^{n}(x+2-n)-2n^{n}\geq 0.$
This can be easily verified by the following facts: $F(n)=F^{\prime}(n)=0$ and $F^{\prime\prime}(x)=\frac{n}{2}(n-1)(n-2)(2n-x)^{n-3}(x-n)>0$ for $n<x<2n$.
@@
Let $X\subset{\mathbb{P}}^{n+1}$ be a Fano hypersurface of degree $1\leq b\leq n$. Then the minimal rational curve is of splitting type ${\mathcal{O}}(1)^{n-b}\oplus{\mathcal{O}}^{b-1}$ so that d=nb+2d=n-b+2 (see [Kol96, Exercise V.4.4]). The volume of $X$ is given by $V_{d}=d^{n}(n-d+2)=(n+2-b)^{n}\cdot b$. One can easily verify that $V_{d}$ is strictly smaller than ${\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$ except when $b=1$ or $b=2$.
Question 4.18.
Assume that a K-semistable Fano manifold $X$ contains a minimal rational curve of degree $3\leq d\leq n-1$ in $X$. Is it true that ${\rm vol}(X)=(-K_{X})^{n}\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$?
In the first version, this was claimed as a result. However due to the issue discussed in Remark 4.4, the question is left un-answered. On the other hand, our current results imply that the answer is affirmative if we instead assume $d\in\{2,n,n+1\}$.
@@
If $a_{0}=\cdots=a_{n-1}=1,\,a_{n}=2,\,a_{n+1}=n+1$ and $d=2n+2$. Namely, $X_{d}$ is a general degree $2n+2$ hypersurface in ${\mathbb{P}}(1^{n},2,n+1)$. Then $(-K_{X_{2n+2}})^{n}=1$. By [IF00, Thm 8.1], we know that $X_{2n+2}$ is smooth when $n$ is even. By [ST24, Thm 1.3], we know $\delta(X_{2n+2})\geqslant(n+1)/2>1$, so $X_{2n+2}$ is a smooth K-stable Fano manifold with minimal anti-canonical volume $(-K_{X_{2n+2}})^{n}=1$ for all even $n$. When $n$ is odd, the general hypersurface $X_{2n+2}\subset{\mathbb{P}}(1^{n},2,n+1)$ is singular. (After communicating with K. Fujita, we find that this type of example already appears in [KSC04, Theorem 5.22].)
@@
Let $Y$ be the double cover of ${\mathbb{P}}^{n}$ ramified along a degree $2n$ smooth hypersurface $D$. Then by Hurwitz’s formula, $K_{Y}=\phi^{*}(K_{{\mathbb{P}}^{n}}+\frac{1}{2}D)$. So $(-K_{Y})^{n}=\deg(\phi)\cdot(-K_{{\mathbb{P}}^{n}}-\frac{1}{2}D)^{n}=2$. It was first shown in [AGP06, Theorem 3.2] that $Y$ admits Kähler-Einstein metric, hence K-stable (See also [Der16] and [AZ22, Corollary 4.9(3)] for a purely algebraic proof of K-stable). Here we illustrate a different proof based on the result of [LZ22, Zhu21], we know K-stability of $Y$ is equivalent to the K-stability of the log Fano pair $({\mathbb{P}}^{n},\frac{1}{2}D)$. Since we assume $D$ is a smooth hypersurface, then the log Fano pair $({\mathbb{P}}^{n},(\frac{n+1}{2n}-\varepsilon)D)$ is klt, so it is uniformly K-stable by [ADL23, Theorem 2.10] for $0<\varepsilon\ll 1$. Since $D\sim_{{\mathbb{Q}}}-K_{{\mathbb{P}}^{n}}$, by the interpolation of K-stability ([ADL24, Prop 2.13]), we conclude that $({\mathbb{P}}^{n},\frac{1}{2}D)$ is K-stable. Thus, $Y$ is a smooth K-stable Fano manifold with volume $(-K_{Y})^{n}=2$ for all dimensions $n\geqslant 2$. Note that $Y$ contains a degree 2 free-immersed rational curve with n1n-1 nodes (see [Kol96, IV.2.12.3]).